### v1

@@ Abstract
Weakly supervised audio-visual video parsing (AVVP) methods aim to detect audible-only, visible-only, and audible-visible events using only video-level labels. Existing approaches tackle this by leveraging unimodal and cross-modal contexts. However, we argue that while cross-modal learning is beneficial for detecting audible-visible events, in the weakly supervised scenario, it negatively impacts unaligned audible or visible events by introducing irrelevant modality information. In this paper, we propose CoLeaF, a novel learning framework that optimizes the integration of cross-modal context in the embedding space such that the network explicitly learns to combine cross-modal information for audible-visible events while filtering them out for unaligned events. Additionally, as videos often involve complex class relationships, modelling them improves performance. However, this introduces extra computational costs into the network. Our framework is designed to leverage cross-class relationships during training without incurring additional computations at inference. Furthermore, we propose new metrics to better evaluate a method’s capabilities in performing AVVP. Our extensive experiments demonstrate that CoLeaF significantly improves the state-of-the-art results by an average of 1.9% and 2.4% F-score on the LLP and UnAV-100 datasets, respectively.
@@ 1 Introduction
To address the weakly supervised AVVP task, all existing approaches, such as [17, 24, 3], incorporate unimodal and cross-modal contexts. However, we argue that while extracting cross-modal information is beneficial for multi-modal learning tasks, in the weakly supervised scenario, it can negatively impact the detection of audible-only and visible-only events (i.e., unaligned audible-visible events) by introducing irrelevant modality information. We investigated this in Table 2 by comparing the performance of the state-of-the-art network CMPAE [3] with and without using cross-modal contexts. The results show that leveraging cross-modal contexts improves the performance on average in detecting audible-visible events, while in detecting unaligned events, a network focusing solely on unimodal information yields better results. These findings motivated us to raise this question: Can we optimize the integration of cross-modal context in the embedding space such that the network explicitly learns to leverage cross-modal information for audible-visible events, while filtering them out for unaligned events? To achieve this, we introduce CoLeaF, a novel learning framework for the weakly supervised AVVP task. CoLeaF relies on two network branches: Reference and Anchor. The Reference branch focuses on unimodal information, while the Anchor branch incorporates unimodal and cross-modal contexts. Both branches process the same audio and visual input signals and are trained simultaneously using the video-level labels. Additionally, during training, the Reference branch is leveraged by a novel contrastive objective to optimize cross-modal learning in the Anchor branch based on the event types present in the input video. The proposed objective encourages Anchor to learn embedding representations similar to those learned in Reference, but the strength of the encouragement adapts to the degree of unaligned audible-visible events present in the input video. More unaligned events lead to a stronger encouragement, while fewer unaligned events lead to a weaker encouragement. The unalignment degree is computed from the Reference’s output predictions. Since the optimization process relies on the performance of Reference, its effectiveness may diminish if Reference is solely trained on ‘modality-agnostic’ video-level labels. To address this without requiring additional annotations, the Anchor branch collaboratively distils free modality-aware pseudo-labels for the Reference branch. After training, the Anchor branch is deployed for inference.
@@
To effectively perform the AVVP task, we should exploit unimodal and cross-modal contexts while aware of class dependencies. For example, this is often seen when musical instruments are played alongside people singing. Authors in [12, 13] design networks to explicitly learn cross-class relationships. However, their approaches are computationally expensive. For example, in [12], to capture cross-class dependencies, they extend their model to map each temporal segment of the video into the maximum number of event classes, and they apply attention layers within the event classes for each segment individually. For a video with a length of T segments and C class categories, this led to a significant computational overhead with a ratio of $T\times C\times C$. This key challenge motivated us to raise this novel question: Can we take advantage of explicitly modelling co-occurrence class relationships without imposing the overhead of heavy computational costs? To achieve this, in our framework, the Reference branch that is employed only during training is designed to explicitly learn co-occurrence class relationships and distil this knowledge to the Anchor branch. This novel design allows us to benefit from cross-class relationships during training without introducing computational overhead on the Anchor branch that is employed at inference.
For a deeper understanding of how our proposed framework models unaligned events, we introduce novel metrics for AVVP. Current metrics evaluate the method’s performance in detecting audible-only and visible-only events by comparing the estimated audio and visual event proposals individually with audio and visual ground-truth labels. In our proposed metrics, we consider both audio and visual modalities, as comparing them individually can lead to the mistaken interpretation of an audible-visible event as both audible-only and visible-only.
@@
Our key contributions are summarized as: (i) we pioneer the first exploration in explicitly optimizing the integration of cross-modal context for the weakly supervised AVVP task, (ii) we propose an approach that allows the network to explicitly learn cross-class relationships without adding computational overhead during inference. This can benefit various computer vision tasks that demand a deeper understanding of interactions between classes, (iii) we introduce novel metrics for effective evaluation of AVVP which are beneficial for future development of this task, (iv) we evaluate CoLeaF on two challenging publicly available datasets and improve the state-of-the-art results on average by 1.9% and 2.4% F-score on LLP and UnAV-100 datasets, respectively, and (v) we perform extensive ablation studies to evaluate our proposed framework’s design.
@@ 3.1 Preliminaries
Figure 3: The overall scheme of our proposed learning framework CoLeaF for the weakly supervised AVVP task. CoLeaF has two network branches, Reference and Anchor. The Reference branch focuses on unimodal information and cross-class relationships, while the Anchor branch exploits both unimodal and cross-modal contexts. The branches are trained simultaneously through the conventional video-level losses $\{\mathcal{{L}}^{Ref}_{video},\mathcal{{L}}^{Anchr}_{video}\}$ and our novel contrastive and knowledge distillation losses $\{\mathcal{L}^{Anch}_{Evt},\mathcal{L}^{Ref}_{SelfMo},\mathcal{L}^{Anchr}_{CoCls}\}$. In inference, Anchor is deployed for the AVVP task.
@@ 3.2 CoLeaF: Contrastive-Collaborative Learning Framework
We take the outputs of self-attention layers which correspond to the audio and visual input tokens, $\ddot{F}^{\phi}=\{\ddot{x}^{\phi}_{n}\}^{T}_{n=1}$, to predict segment-level audio and visual event probabilities through a fully connected layer as $\{\ddot{p}^{\phi}_{t}\}^{T}_{t=1}=Sigmoid(FC_{\phi}(\ddot{F}^{\phi}))$, where $\ddot{p}^{\phi}_{t}\in\rm{I\!R}^{1\times C}$. Subsequently, the audio and visual event probabilities are computed as
@@ 5.3 Generalization on the Large-Scale UnAV-100 Dataset
To demonstrate the superior generalization of our proposed framework on a different, yet challenging dataset, we present its results to the large-scale UnAV-100 dataset, a benchmark for the Dense Audio-Visual Event Localization (DAVE) task [4]. UnAV-100 includes videos with multiple events and encompasses a diverse range of 100 distinct audio-visible action types, four times more than the LLP dataset. Furthermore, UnAV-100 features untrimmed videos ranging from 0.2 to 60 seconds, significantly increasing the dataset’s complexity. To obtain weakly supervised results, we derived video-level labels from the videos’ temporal annotations. The results of other state-of-the-art approaches were provided through their available code. Comparative results on UnAV-100 using AV F-score are available in Table 5. It is shown that CoLeaF outperforms the other approaches by 1.8% and 3.0% AV F-score at the segment and event levels, respectively, and achieves state-of-the-art results on this dataset.

### v4

@@ Abstract
Weakly supervised audio-visual video parsing (AVVP) methods aim to detect audible-only, visible-only, and audible-visible events using only video-level labels. Existing approaches tackle this by leveraging unimodal and cross-modal contexts. However, we argue that while cross-modal learning is beneficial for detecting audible-visible events, in the weakly supervised scenario, it negatively impacts unaligned audible or visible events by introducing irrelevant modality information. In this paper, we propose CoLeaF, a novel learning framework that optimizes the integration of cross-modal context in the embedding space such that the network explicitly learns to combine cross-modal information for audible-visible events while filtering them out for unaligned events. Additionally, as videos often involve complex class relationships, modelling them improves performance. However, this introduces extra computational costs into the network. Our framework is designed to leverage cross-class relationships during training without incurring additional computations at inference. Furthermore, we propose new metrics to better evaluate a method’s capabilities in performing AVVP. Our extensive experiments demonstrate that CoLeaF significantly improves the state-of-the-art results by an average of 1.9% and 2.4% F-score on the LLP and UnAV-100 datasets, respectively. Code is available at: https://github.com/faeghehsardari/coleaf.
@@ 1 Introduction
To address the weakly supervised AVVP task, all existing approaches, such as [17, 24, 3], incorporate unimodal and cross-modal contexts. However, we argue that while extracting cross-modal information is beneficial for multi-modal learning tasks, in the weakly supervised scenario, it can negatively impact the detection of audible-only and visible-only events (i.e., unaligned audible-visible events) by introducing irrelevant modality information. We investigated this in Table 2 by comparing the performance of CMPAE [3] with and without using cross-modal contexts. The results show that leveraging cross-modal contexts improves the performance on average in detecting audible-visible events, while in detecting unaligned events, a network focusing solely on unimodal information yields better results. These findings motivated us to raise this question: Can we optimize the integration of cross-modal context in the embedding space such that the network explicitly learns to leverage cross-modal information for audible-visible events, while filtering them out for unaligned events? To achieve this, we introduce CoLeaF, a novel learning framework for the weakly supervised AVVP task. CoLeaF relies on two network branches: Reference and Anchor. Reference focuses on unimodal information, while Anchor incorporates unimodal and cross-modal contexts. Both branches process the same audio and visual input signals and are trained simultaneously using the video-level labels. Additionally, during training, Reference is leveraged by a novel contrastive objective to optimize cross-modal learning in Anchor based on the event types present in the input video. The proposed objective encourages Anchor to learn embedding representations similar to those learned in Reference, but the strength of the encouragement adapts to the degree of unaligned audible-visible events present in the input video. More unaligned events lead to a stronger encouragement, while fewer unaligned events lead to a weaker encouragement. The unalignment degree is computed from the Reference’s output predictions. Since the optimization process relies on the performance of Reference, its effectiveness may diminish if Reference is solely trained on ‘modality-agnostic’ video-level labels. To address this without requiring additional annotations, Anchor collaboratively distils free modality-aware pseudo-labels for Reference. After training, Anchor is deployed for inference.
@@
To effectively perform the AVVP task, we should exploit unimodal and cross-modal contexts while aware of class dependencies. For example, this is often seen when musical instruments are played alongside people singing. Authors in [12, 13] design networks to explicitly learn cross-class relationships. However, their approaches are computationally expensive. For example, in [12], to capture cross-class dependencies, they extend their model to map each temporal segment of the video into the maximum number of event classes, and apply attention layers within the event classes for each segment individually. For a video with a length of T segments and C class categories, this led to a significant computational overhead with a ratio of $T\times C\times C$. This key challenge motivated us to raise this novel question: Can we take advantage of explicitly modelling co-occurrence class relationships without imposing the overhead of heavy computational costs? To achieve this, in CoLeaF, Reference that is employed only during training is designed to explicitly learn co-occurrence class relationships and distil this knowledge to Anchor. This novel design allows us to benefit from cross-class relationships during training without introducing computational overhead on Anchor that is employed at inference.
@@
For a deeper understanding of how our proposed framework models unaligned events, we introduce novel metrics for AVVP. Current metrics evaluate the method’s performance in detecting audible-only and visible-only events by comparing the estimated audio and visual event proposals individually with audio and visual ground-truth labels. In our proposed metrics, we consider both audio and visual modalities, as comparing them individually can lead to the mistaken interpretation of an audible-visible event as both audible-only and visible-only. Our key contributions are summarized as: (i) we pioneer the first exploration in explicitly optimizing the integration of cross-modal context for the weakly supervised AVVP task, (ii) we propose an approach that allows the network to explicitly learn cross-class relationships without adding computational overhead during inference. This can benefit various computer vision tasks that demand a deeper understanding of interactions between classes, (iii) we introduce novel metrics for effective evaluation of AVVP which are beneficial for future development of this task, (iv) we evaluate CoLeaF on two challenging publicly available datasets and improve the state-of-the-art results on average by 1.9% and 2.4% F-score on LLP and UnAV-100 datasets, respectively, and (v) we perform extensive ablation studies to evaluate our proposed framework’s design.
@@ 3.1 Preliminaries
Figure 3: The overall scheme of CoLeaF for the weakly supervised AVVP task. CoLeaF has two network branches, Reference and Anchor. The Reference branch focuses on unimodal information and cross-class relationships, while the Anchor branch exploits both unimodal and cross-modal contexts. The branches are trained simultaneously through the conventional video-level losses $\{\mathcal{{L}}^{Ref}_{video},\mathcal{{L}}^{Anchr}_{video}\}$ and our novel contrastive and knowledge distillation losses $\{\mathcal{L}^{Anch}_{Evt},\mathcal{L}^{Ref}_{SelfMo},\mathcal{L}^{Anchr}_{CoCls}\}$. In inference, Anchor is deployed for the AVVP task.
@@ 3.2 CoLeaF: Contrastive-Collaborative Learning Framework
We take the outputs of self-attention layers which correspond to the audio and visual input tokens, $\ddot{F}^{\phi}=\{\ddot{x}^{\phi}_{n}\}^{T}_{n=1}$, to predict segment-level audio and visual event probabilities through a fully connected layer as $\{\ddot{p}^{\phi}_{t}\}^{T}_{t=1}=Sigmoid(FC_{\phi}(\ddot{F}^{\phi}))$, where $\ddot{p}^{\phi}_{t}\in\rm{I\!R}^{1\times C}$. Subsequently, the audio and visual event probabilities are computed as ${\ddot{\mathcal{P}}}^{\phi}=\sum_{t=1}^{T}(\ddot{w}^{\phi}_{t}\odot\ddot{p}^{\phi}_{t})$
@@ 5.3 Generalization on the Large-Scale UnAV-100 Dataset
To demonstrate the superior generalization of our proposed framework on a different, yet challenging dataset, we present its results to the large-scale UnAV-100 dataset, a benchmark for the Dense Audio-Visual Event Localization (DAVE) task [4]. UnAV-100 includes videos with multiple events and encompasses a diverse range of 100 distinct audio-visible action types, four times more than the LLP dataset. Furthermore, UnAV-100 features untrimmed videos ranging from 0.2 to 60 seconds, significantly increasing the dataset’s complexity. To obtain weakly supervised results, we derived video-level labels from the videos’ temporal annotations (see supp. file for the implementation details for this task). The results of other state-of-the-art approaches were provided through their available code. Comparative results on UnAV-100 using AV F-score are available in Table 5. It is shown that CoLeaF outperforms the other approaches by 1.8% and 3.0% AV F-score at the segment and event levels, respectively, and achieves state-of-the-art results on this dataset.
@@ 6 Conclusion
Acknowledgement
This research is supported by UKRI EPSRC Platform Grant EP/P022529/1, and EPSRC BBC Prosperity Partnership AI4ME: Future Personalised Object-Based Media Experiences Delivered at Scale Anywhere EP/V038087/1.